# Leveraging Hamiltonian Simulation Techniques to Compile Operations on Bosonic Devices

### v2

@@ Abstract
Circuit quantum electrodynamics enables the combined use of qubits and oscillator modes. Despite a variety of available gate sets, many hybrid qubit-boson (i.e., qubit-oscillator) operations are realizable only through optimal control theory, which is oftentimes intractable and uninterpretable. We introduce an analytic approach with rigorously proven error bounds for realizing specific classes of operations via two matrix product formulas commonly used in Hamiltonian simulation, the Lie–Trotter and Baker–Campbell–Hausdorff product formulas. We show how this technique can be used to realize a number of operations of interest, including polynomials of annihilation and creation operators, namely $(a)^{p}(a^{\dagger})^{q}$ for integer $p,q$. We show examples of this paradigm including obtaining universal control within a subspace of the entire Fock space of an oscillator, state preparation of a fixed photon number in the cavity, simulation of the Jaynes–Cummings Hamiltonian, and simulation of the Hong-Ou-Mandel effect. This work demonstrates how techniques from Hamiltonian simulation can be applied to better control hybrid qubit-boson devices.
@@ 1 Introduction
Today, many quantum computing architectures are homogeneous, with the same type of qubit used throughout the device. From devices made of superconducting qubits [arute2019quantum, dial2022moving, reagor2018demonstration] to ion trap qubits [wright2019benchmarking], prior work largely focuses on linking qubits of the same type together in fault-tolerant ways. However, there is emerging work [C2QA_ISA, C2QA_LGT, chakram2021seamless, stavenger2022bosonic, stein2023microarchitectures] studying the use of heterogeneous quantum computers that leverage two or more types of quantum architectures (e.g., qubits and oscillator modes). Heterogeneous devices hold promise because they can be tailored for specific physical simulation problems, which would be especially useful in applications like material discovery [Holstein_Knorzer_2022], molecular simulation [Wang2020FCFs, WangConicalIntersection], topological models [PhysRevB.98.174505] or lattice gauge theories [C2QA_LGT].
@@
We develop two parallel approaches, one which primarily uses the creation and destruction operators which we refer to as based on ‘Fock methods,’ and the other primarily relying on position and momentum operators which we refer to as based on ‘phase-space methods.’ We demonstrate that both methods can be used to generate an ISA for qumode-qubit devices. Our methods obtain almost-linear asymptotic scaling.
@@ 2 Preliminaries
We review the mathematical properties of qumode quantum mechanics that are needed to understand the basic operations considered for the ISA architecture that we consider. Specifically, we present an analytic instruction set architecture (ISA) based on the Lie-Trotter-Suzuki (Trotter) and Baker-Campbell-Hausdorff (BCH) decompositions for decomposition of gates of the form $U=e^{i\hat{H}\sigma^{j}}$, where $\hat{H}$ is a Hermitian operator composed of phase-space operators and Pauli gates. Before jumping into the specific details of our gate operations, we need to review the basics of qumode quantum mechanics as well as the mathematical results needed to use these qumode operations to compile a given unitary transformation.
@@ 2.1.1 Representing the qumode
Truncation of the Hilbert space is required to provide error bounds, otherwise the remainder terms become undefined. Provided that an appropriate cutoff is picked for the system, the discrepancies between the truncated and untruncated systems will often be negligible. For notational clarity, we assume a cutoff of $\Lambda$ for all further equations and assume the annihilation and creation operators implicitly have the projectors $P_{\Lambda}$.
@@ 2.1.2 Representing operations and measurements via block encodings
    Qubit-exclusive: these include typical qubit gates like the Pauli ($X,Y,Z$) gates, Hadamard ($H$) gate
    $H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},$ (7)
    and phase ($S$) gate
    $S=\begin{bmatrix}1&0\\ 0&i\end{bmatrix}.$ (8)
@@
    Qumode-qubit entangling: there are several entangling operations between individual qumodes and qubits that widely appear in the circuit QED literature. Two common operations we consider are the controlled displacement operation [eickbusch2021fast] $e^{-i\sigma^{z}\otimes(\alpha a^{\dagger}+\alpha a)}$ and the Selective Number-dependent Arbitrary Phase (SNAP) gate [SNAP-PhysRevLett.115.137002] $e^{-i\sigma^{z}\otimes\sum_{n}\alpha_{n}\hat{P}_{n}}$, where $\hat{P}_{n}=|n\rangle\langle n|$ is the projector onto the $n$th Fock state.
@@ 4.3 State preparation from the vacuum
Figure 4: Unprotected $T_{2}$ Gate: plots visualize the unitary’s elements. Brighter colors represent higher absolute values / magnitudes of the matrix entries. (a) is the exact form of the $T_{2}$ gate with selected $t$ and (b) is the BCH-synthesized form. The state |$\ket{j}_{q}\ket{k}_{m}$ has index $j\cdot\Lambda+k$ where $\Lambda$ is the cutoff. By observation, the analytically realized form is accurate.
Figure 5: Protected $T_{2}$ Gate: (a) is the exact form of the protected $T_{2}$ gate with selected $t$ and (b) is the BCH-synthesized form. The state |$\ket{j}_{q}\ket{k}_{m}$ has index $j\cdot\Lambda+k$ where $\Lambda$ is the cutoff. By observation, the analytically realized form is still accurate, albeit with more incurred Trotter error.
@@ Acknowledgements
This project was primarily supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage under contract number DE-SC0012704, (Basic Energy Sciences, PNNL FWP 76274). Eleanor Crane was supported by UCL Faculty of Engineering Sciences and the Yale-UCL exchange scholarship from RIGE (Research, Innovation and Global Engagement), the Princeton MURI award SUB0000082, the DoE QSA, NSF QLCI (award No.OMA-2120757), DoE ASCR Accelerated Research in Quantum Computing program (award No.DESC0020312), and NSF PFCQC program. Micheline B. Soley was supported by the Yale Quantum Institute Postdoctoral Fellowship and the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin-Madison with funding from the Wisconsin Alumni Research Foundation.
@@ Appendix A Obtaining $S_{1}$
The controlled displacement operator with magnitude $\alpha$ is written:

### v3

@@ Abstract
Circuit quantum electrodynamics enables the combined use of qubits and oscillator modes. Despite a variety of available gate sets, many hybrid qubit-boson (i.e., qubit-oscillator) operations are realizable only through optimal control theory, which is oftentimes intractable and uninterpretable. We introduce an analytic approach with rigorously proven error bounds for realizing specific classes of operations via two matrix product formulas commonly used in Hamiltonian simulation, the Lie–Trotter–Suzuki and Baker–Campbell–Hausdorff product formulas. We show how this technique can be used to realize a number of operations of interest, including polynomials of annihilation and creation operators, namely $(a)^{p}(a^{\dagger})^{q}$ for integer $p,q$. We show examples of this paradigm including obtaining universal control within a subspace of the entire Fock space of an oscillator, state preparation of a fixed photon number in the cavity, simulation of the Jaynes–Cummings Hamiltonian, and simulation of the Hong-Ou-Mandel effect. This work demonstrates how techniques from Hamiltonian simulation can be applied to better control hybrid qubit-boson devices.
@@ 1 Introduction
Today, many quantum computing architectures are homogeneous, with the same type of qubit used throughout the device. From devices made of superconducting qubits [arute2019quantum, bravyi2022future, reagor2018demonstration] to ion trap qubits [wright2019benchmarking], prior work largely focuses on linking qubits of the same type together in fault-tolerant ways. However, there is emerging work [C2QA_ISA, crane2024hybrid, chakram2021seamless, stavenger2022bosonic, stein2023microarchitectures] studying the use of heterogeneous quantum computers that leverage two or more types of quantum architectures (e.g., qubits and oscillator modes). Heterogeneous devices hold promise because they can be tailored for specific physical simulation problems, which would be especially useful in applications like material discovery [Holstein_Knorzer_2022], molecular simulation [Wang2020FCFs, WangConicalIntersection], topological models [PhysRevB.98.174505] or lattice gauge theories [C2QA_LGT].
@@
We develop two parallel approaches, one which primarily uses the creation and annihilation operators which we refer to as based on ‘Fock methods,’ and the other primarily relying on position and momentum operators which we refer to as based on ‘phase-space methods.’ We demonstrate that both methods can be used to generate an ISA for qumode-qubit devices. Because our methods operate in both Fock and phase-space, we can achieve transformations that are natively described in either picture; for example, exponentials of polynomials of annihilation and creation operators for Hamiltonian simulation and state preparation in Fock space and controlled parity and beam splitter operators in phase space. Our methods obtain almost-linear asymptotic scaling.
@@ 2 Preliminaries
We review the mathematical properties of qumode quantum mechanics that are needed to understand the basic operations considered for the ISA architecture that we consider. Specifically, we present an analytic instruction set architecture (ISA) based on the Lie–Trotter–Suzuki (Trotter) and Baker–Campbell–Hausdorff (BCH) decompositions for decomposition of gates of the form $U=e^{i\hat{H}\sigma^{j}}$, where $\hat{H}$ is a Hermitian operator composed of phase-space operators and Pauli gates. Before jumping into the specific details of our gate operations, we need to review the basics of qumode quantum mechanics as well as the mathematical results needed to use these qumode operations to compile a given unitary transformation.
@@ 2.1.1 Representing the qumode
Observe that this Hilbert space has dimension $\Gamma+1$ corresponding to the Fock states from |$\ket{0}$ to |$\ket{\Gamma}$. Truncation of the Hilbert space is required to provide error bounds, otherwise the remainder terms become undefined. Provided that an appropriate cutoff is picked for the system, the discrepancies between the truncated and untruncated systems will often be negligible. For notational clarity, we assume a cutoff of $\Lambda$ for all further equations and assume the annihilation and creation operators implicitly have the projectors $P_{\Lambda}$.
@@ 2.1.2 Representing operations and measurements via block encodings
    Qubit-exclusive: these include typical qubit gates like the Pauli ($X,Y,Z$) gates, ADDED below:
    $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\qquad Y=\begin{bmatrix}0&-i\\ i&0\end{bmatrix},\qquad\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},$ (7)
    and Hadamard ($H$) and phase ($S$) gates
    $H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\qquad S=\begin{bmatrix}1&0\\ 0&i\end{bmatrix}.$ (8)
@@
    Qumode-qubit entangling: there are several entangling operations between individual qumodes and qubits that widely appear in the circuit QED literature. Two common operations we consider are the conditional displacement operation [eickbusch2021fast] $e^{-i\sigma^{z}\otimes(\alpha a^{\dagger}+\alpha a)}$ and the Selective Number-dependent Arbitrary Phase (SNAP) gate [SNAP-PhysRevLett.115.137002] $e^{-i\sigma^{z}\otimes\sum_{n}\alpha_{n}\hat{P}_{n}}$, where $\hat{P}_{n}=|n\rangle\langle n|$ is the projector onto the $n$th Fock state.
@@ 4.3 State preparation from the vacuum
As can be seen in Figure 6, the protected gate has higher compilation errors than the unprotected gate. This is likely due to the fact that the protected formula requires an additional application of the Trotter product formula. This will further incur commutator error over the unprotected gate.
Figure 4: Unprotected $T_{2}$ Gate: plots visualize the unitary’s elements. Brighter colors represent higher absolute values / magnitudes of the matrix entries. Left is the exact form of the $T_{2}$ gate with selected $t$ and right is the BCH-synthesized form. The state |$\ket{j}_{q}\ket{k}_{m}$ has index $j\cdot\Lambda+k$ where $\Lambda$ is the cutoff. By observation, the analytically realized form is accurate.
Figure 5: Protected $T_{2}$ Gate: left is the exact form of the protected $T_{2}$ gate with selected $t$ and right is the BCH-synthesized form. The state |$\ket{j}_{q}\ket{k}_{m}$ has index $j\cdot\Lambda+k$ where $\Lambda$ is the cutoff. By observation, the analytically realized form is still accurate, albeit with more incurred Trotter error.
Figure 6: Absolute element-wise error when compiling $T_{2}$ gate via scheme: left is when the compilation target is the unprotected gate, while right is when the target is the protected gate. The protected gate has higher synthesis errors.
@@ 5 Discussion
Author Contributions
SMG and NW conceived of the instruction set architecture and applications. CK derived and obtained analytic error bounds, developed the state preparation subroutine, and mathematically formalized the scheme. MBS developed compilation strategies for practical implementation, investigated the efficacy of the approach for physical applications, and performed numerical simulations of gate compilation and quantum dynamics. EC advocated for clarified exposition across all applications. All authors contributed to the manuscript, with EC, CK, and MBS producing the initial and revised manuscripts.
@@ Acknowledgements
This project was primarily supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage under contract number DE-SC0012704, (Basic Energy Sciences, PNNL FWP 76274). C2QA led this research. Christopher Kang was also funded in part by the STAQ project under award NSF Phy-232580; in part by the US Department of Energy Office of Advanced Scientific Computing Research, Accelerated Research for Quantum Computing Program; and in part by the NSF Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks (NSF Award 2016136), in part based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, and in part by the Army Research Office under Grant Number W911NF-23-1-0077. Micheline B. Soley was supported by the Yale Quantum Institute Postdoctoral Fellowship and the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin-Madison with funding from the Wisconsin Alumni Research Foundation. Eleanor Crane was supported by UCL Faculty of Engineering Sciences and the Yale-UCL exchange scholarship from RIGE (Research, Innovation and Global Engagement), the Princeton MURI award SUB0000082, the DoE QSA, NSF QLCI (award No.OMA-2120757), DoE ASCR Accelerated Research in Quantum Computing program (award No.DESC0020312), and NSF PFCQC program.
The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Obtaining $S_{1}$
The conditional displacement operator with magnitude $\alpha$ is written: